jay is 12 years older than mario. If mario's age 2 years from now is half jay's age six years from now, how ild will each be in 10 years?
Answer by mananth(16949)

jay is 12 years older than mario. If mario's age 2 years from now is half jay's age six years from now, how ild will each be in 10 years?
...
let mario be x
Jay = x+12
...
2 years from now
Mario = x+2
6 years from now Jay's age
x+12+6
=x+18
...
x+2 = 1/2 (x+18)
2(x+2)=x+18
2x+4 = x+18
2x-x=18-4
x= 14 Mario's age now
x+12 = 14+12 = 26 years Jay's age
..
10 years later
Mario = 24 years
Jay = 36 years